# Implicit Regularization of Infinitesimally-perturbed Gradient Descent Toward Low-dimensional Solutions

### v1

@@ 3 Establishing Convergence of IPGD to an Approximate SOSP
where we denote $\lambda_{\min}:=\lambda_{\min}(\nabla^{2}f(\mathbf{z}))\leq-\sqrt{\rho\epsilon}$. Next, we use induction to show that $q_{t}\leq\frac{1}{2}p_{t}$ for all $0\leq t\leq T$. Indeed, the base case for $t=0$ holds since $q_{0}=0$. Now, suppose that $q_{s}\leq\frac{1}{2}p_{s}$ for all $10\leq s\leq t-$ 1. Then, an upper bound on $q_{t}$ can be derived as follows
$q_{t}$ $\leq\eta\sum_{s=0}^{t-1}\left\lVert(\mathbf{I}-\eta\mathbf{H})^{t-s-1}\right\rVert\left\lVert\bm{\Delta}_{s}\right\rVert\left\lVert\bm{\delta}_{s}\right\rVert$ (33)
$\leq\eta\sum_{s=0}^{t-1}(1-\eta\lambda_{\min})^{t-s-1}\left\lVert\bm{\Delta}_{s}\right\rVert\left\lVert\bm{\delta}_{s}\right\rVert$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\rho\phi\cdot\eta\sum_{s=0}^{t-1}(1-\eta\lambda_{\min})^{t-s-1}\cdot\frac{3}{2}p_{s}$
$\leq\frac{3}{2}\eta\rho\phi\cdot\sum_{s=0}^{t-1}(1-\eta\lambda_{\min})^{t-s-1}\cdot(1-\eta\lambda_{\min})^{s}p_{0}$
$=\frac{3}{2}\eta t\rho\phi\cdot p_{t-1}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{1}{2}p_{t}.$
Here in $(a)$, we use $\left\lVert\bm{\Delta}_{t}\right\rVert\leq\rho\phi$ and the assumption $\left\lVert\bm{\delta}_{s}\right\rVert\leq p_{s}+q_{s}\leq\frac{3}{2}p_{s}$ for all $10\leq s\leq t-$ 1. In $(b)$, we use the fact that $t\leq\frac{2}{\eta\sqrt{\rho\epsilon}}\log\left(\frac{1}{\omega}\right)$ and $\phi=\frac{\sqrt{\epsilon}}{6\sqrt{\rho}}\log^{-1}\left(\frac{1}{\omega}\right)\leq\frac{1}{4}$, which implies
@@
Therefore, within $T=\log\left(\frac{4\phi}{\omega}\right)/\log\left(1+\eta\sqrt{\rho\epsilon}\right)$ iterations, we must have $\left\lVert\bm{\delta}_{T}\right\rVert\geq 2\phi$. Note that $\phi\leq\frac{1}{4}$ and $\log(1+\eta\sqrt{\rho\epsilon})\geq\frac{\eta\sqrt{\rho\epsilon}}{1+\eta\sqrt{\rho\epsilon}}\geq\frac{1}{2}\eta\sqrt{\rho\epsilon}$. Hence, we have $T\leq\frac{2}{\eta\sqrt{\rho\epsilon}}\log\left(\frac{1}{\omega}\right)$. This further implies
@@
$\mathcal{X}_{\mathrm{stuck}}:=\left\{\mathbf{x}_{0}\in\mathbb{B}_{\mathbf{z}}(\gamma):\right.$ $f(\mathbf{x}_{T})-f(\mathbf{z})\geq-F$ (40)
[…]
@@
Using Lemma 14, we show that for any two GD sequences with initial points $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}\in\mathbb{B}_{\mathbf{z}}(\gamma)$ satisfying $\mathbf{x}_{0}-\mathbf{x}_{0}^{\prime}=\omega\mathbf{v}$, at least one of them lies outside the stuck region. To this goal, we write
@@
$\mathbb{P}\left(\mathbf{x}_{0}\in\mathcal{X}_{\mathrm{stuck}}\right)=\frac{\operatorname{vol}\left(\mathcal{X}_{\text{stuck}}\right)}{\operatorname{vol}\left(\mathbb{B}^{d}(\gamma)\right)}\leq\frac{\omega\cdot\operatorname{vol}\left(\mathbb{B}^{d-1}(\gamma)\right)}{\operatorname{vol}\left(\mathbb{B}^{d}(\gamma)\right)}=\frac{\omega\Gamma(d/2+1)}{\sqrt{\pi}\gamma\Gamma(d/2+1/2)}$ $\leq\frac{\omega}{\gamma}\sqrt{\frac{d+1}{2\pi}}$ (43)
$\leq\frac{\chi}{N_{\mathrm{perturb}}}.$
Here, $\operatorname{vol}(\mathcal{X})$ denotes the volume of the set $\mathcal{X}$. For the $d$-dimensional unit ball, we have $\operatorname{vol}(\mathbb{B}^{d}(1))=\frac{\pi^{d/2}}{\Gamma\left(\frac{d}{2}+1\right)}$, where $\Gamma(\cdot)$ is the Gamma function defined as $\Gamma\left(\frac{d}{2}+1\right)=\sqrt{\pi}\left(d-\frac{1}{2}\right)\cdot\left(d-\frac{3}{2}\right)\cdot\ldots\cdot\frac{1}{2}$ for $d=1,2,\dots$. Moreover, in the second inequality, we use $\frac{\Gamma(x+1)}{\Gamma(x+1/2)}<\sqrt{x+\frac{1}{2}}$ for all $x\geq 0$. ∎
@@ 4 Establishing Implicit Regularization of IPGD
Upon defining $\mathbf{u}^{\star}=\operatorname*{arg\,min}_{\mathbf{u}\in\mathcal{M}}\left\lVert\mathbf{u}-(\mathbf{x}+\mathbf{y})\right\rVert$ and $\mathbf{v}^{\star}=\operatorname*{arg\,min}_{\mathbf{v}\in\mathcal{M}}\left\lVert\mathbf{v}-\mathbf{x}\right\rVert$, we have:
@@
$\left\lVert\left(\mathbf{I}-\eta\nabla^{2}f\big{(}\bm{\mathbf{x}}^{\sharp}_{t}\big{)}\right)\mathbf{x}_{t}^{\perp}\right\rVert^{2}$ $=\left\lVert\sum_{i=1}^{d}(1-\eta\lambda_{i})\left\langle\mathbf{v}_{i},\bm{\mathbf{x}}^{\perp}_{t}\right\rangle\mathbf{v}_{i}\right\rVert^{2}$ (50)
$=\sum_{i=1}^{d}(1-\eta\lambda_{i})^{2}\left\langle\mathbf{v}_{i},\bm{\mathbf{x}}^{\perp}_{t}\right\rangle^{2}$
@@
$\left\lVert\left(\mathbf{I}-\eta\nabla^{2}f\big{(}\bm{\mathbf{x}}^{\sharp}_{t}\big{)}\right)\mathbf{x}_{t}^{\perp}\right\rVert^{2}$ (52)
$\leq\sum_{i:\lambda_{i}>0}\left(1-\eta\lambda_{i}\right)\left\langle\mathbf{v}_{i},\bm{\mathbf{x}}^{\perp}_{t}\right\rangle^{2}+\sum_{i:\lambda_{i}<0}\left(1-3\eta\lambda_{i}\right)\left\langle\mathbf{v}_{i},\bm{\mathbf{x}}^{\perp}_{t}\right\rangle^{2}$
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert^{2}+\eta\left(-\left\langle\bm{\mathbf{x}}^{\perp}_{t},\nabla^{2}_{+}f\big{(}\bm{\mathbf{x}}^{\sharp}_{t}\big{)}\bm{\mathbf{x}}^{\perp}_{t}\right\rangle-3\left\langle\bm{\mathbf{x}}^{\perp}_{t},\nabla^{2}_{-}f\big{(}\bm{\mathbf{x}}^{\sharp}_{t}\big{)}\bm{\mathbf{x}}^{\perp}_{t}\right\rangle\right)$
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert^{2}+\eta\left(r_{-}(\mathbf{x}_{t})+3r_{+}(\mathbf{x}_{t})\right)\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert^{2}$
=$\stackrel{{\scriptstyle(c)}}{{=}}(1+\eta r(\mathbf{x}_{t}))\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert^{2}.$
@@
$\geq\left(1+\frac{\eta}{2}R(\tau)+\frac{\eta\rho}{2}b_{t_{N_{\mathrm{perturb}}}-1}\right)b_{t_{N_{\mathrm{perturb}}}-1}+\gamma$
$=b_{t_{N_{\mathrm{perturb}}}}+\gamma.$
@@
$\geq\left(1+\frac{\eta}{2}R(\tau)+\frac{\eta\rho}{2}\left(b_{t_{N_{\mathrm{perturb}}}}+\gamma\right)\right)\left(b_{t_{N_{\mathrm{perturb}}}}+\gamma\right)$
$\geq b_{t_{N_{\mathrm{perturb}}}+1}.$
@@
Note that $a_{t+1}=a_{0}\prod_{s=0}^{t}\left(1+\eta R(\tau)+\eta\rho a_{s}\right)$. Hence, it suffices to prove $a_{t}\leq\frac{1}{\eta\rho T}$ for all $10\leq t\leq T-$ 1. We prove this by induction on $t$. It is easy to verify the base case $t=0$. Now, assume that $a_{s}\leq\frac{1}{\eta\rho T}$ for all $0\leq s\leq t$. Then, we have
$a_{t+1}$ $=a_{0}\prod_{s=0}^{t}\left(1+\eta R(\tau)+\eta\rho a_{s}\right)$ (59)
$\leq a_{0}\left(1+\eta R(\tau)+\frac{1}{T}\right)^{t}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}a_{0}\exp\left\{\eta R(\tau)t+\frac{t}{T}\right\}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}10a_{0}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\frac{1}{\eta\rho T},$
@@
First, according to Lemmas 18 and 19, we have that for all $0\leq t\leq T$,
@@
where the final inequality uses the facts that $\left\lVert\mathbf{x}_{0}^{\perp}\right\rVert\leq\frac{\epsilon}{10L}$, $N_{\mathrm{perturb}}=\frac{C\sqrt{\rho}\Delta_{f}\left(\log^{3}\left(\frac{1}{\gamma}\right)+\log^{3}\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)}{\epsilon^{3/2}}$, $\gamma=O\left(\frac{\epsilon^{5/2}}{\sqrt{\rho}L\Delta_{f}}\log^{-3}\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)$, and $R(\tau)=O(\frac{1}{\eta T})$. Given the fact $\left\lVert\mathbf{x}_{t}^{\perp}\right\rVert\leq\frac{\epsilon}{L}$ for all $0\leq t\leq T$, the gradient and Hessian of the objective function remain Lipschitz continuous along the entire trajectory $\{\mathbf{x}_{t}\}_{t=0}^{T}$. Thus, we can apply Theorem 2, which guarantees that, under the stated conditions, with probability at least $1-\chi$, IPGD outputs a point $\mathbf{x}_{T}$ that is an $\epsilon$-SOSP after $T=O\left(\frac{\Delta_{f}}{\eta\epsilon^{2}}\left(\log^{4}\left(\frac{1}{\gamma}\right)+\log^{4}\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)\right)$ iterations. Combining the above two results ensures that $\mathbf{x}_{T}$ is an $\epsilon$$-\mathcal{M}$-SOSP. ∎
@@ 5 Establishing Improved Convergence of IPGD with Additional Structures
[…]
@@
[…]
$(\textup{I})^{2}$ $=\min_{\mathbf{x}^{\star}\in\mathcal{X}_{\mathcal{M}}}\left\lVert\mathbf{x}_{t}^{\sharp}-\eta\nabla f\big{(}\mathbf{x}_{t}^{\sharp}\big{)}-\mathbf{x}^{\star}\right\rVert^{2}$ (68)
$\leq\left\lVert\mathbf{x}_{t}^{\sharp}-\eta\nabla f\big{(}\mathbf{x}_{t}^{\sharp}\big{)}-\mathcal{P}_{\mathcal{X}_{\mathcal{M}}}\big{(}\mathbf{x}_{t}^{\sharp}\big{)}\right\rVert^{2}$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left(1-\frac{\eta\alpha}{2}\right)\left\lVert\mathbf{x}_{t}^{\sharp}-\mathcal{P}_{\mathcal{X}_{\mathcal{M}}}\big{(}\mathbf{x}_{t}^{\sharp}\big{)}\right\rVert^{2}+\left(\eta^{2}-\frac{\eta}{2\beta}\right)\left\lVert\nabla f\big{(}\mathbf{x}_{t}^{\sharp}\big{)}\right\rVert^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\left(1-\frac{\eta\alpha}{2}\right)\left\lVert\mathbf{x}_{t}^{\sharp}-\mathcal{P}_{\mathcal{X}_{\mathcal{M}}}\big{(}\mathbf{x}_{t}^{\sharp}\big{)}\right\rVert^{2},$
@@
[…]
@@
Lemma 21 (Adapted from Lemma 47 in [53]).
@@
$\leq\left(1-\frac{\eta\alpha}{4}\right)\zeta+4\max_{0\leq t\leq T}\left\lVert\mathbf{x}_{t}^{\perp}\right\rVert$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left(1-\frac{\eta\alpha}{4}\right)\zeta+O\left(\frac{\eta\alpha}{L}\epsilon\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\zeta.$
In $(a)$, we use Equation 73. Step $(b)$ follows from the assumption that $\zeta\geq\frac{\epsilon}{2L}$. This completes the inductive step. Given that we established $\mathbf{x}^{\sharp}_{t}\in\mathcal{N}_{\mathcal{X}_{\mathcal{M}}}(\zeta),0\leq t\leq T$, we can invoke Lemma 20 for any $10\leq t\leq T-$ 1, which yields
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left(1-\frac{\eta\alpha}{4}\right)^{T}\cdot 2\zeta+\frac{8+4\eta L}{\eta\alpha}\max_{0\leq t\leq T}\left\{\left\lVert\mathbf{x}_{t}^{\perp}\right\rVert\right\}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\left(1-\frac{\eta\alpha}{4}\right)^{T}\cdot 2\zeta+\frac{\epsilon}{2L}.$
@@
The proof readily follows by combining Proposition 6 and Proposition 7. ∎

### v2

@@ 3 Establishing Convergence of IPGD to an Approximate SOSP
where we denote $\lambda_{\min}:=\lambda_{\min}(\nabla^{2}f(\mathbf{z}))\leq-\sqrt{\rho\epsilon}$. Next, we use induction to show that $q_{t}\leq\frac{1}{2}p_{t}$ for all $0\leq t\leq T$. Indeed, the base case for $t=0$ holds since $q_{0}=0$. Now, suppose that $q_{s}\leq\frac{1}{2}p_{s}$ for all $0\leq s\leq t-1$. Then, an upper bound on $q_{t}$ can be derived as follows
$q_{t}$ $\leq\eta\sum_{s=0}^{t-1}\left\lVert(\mathbf{I}-\eta\mathbf{H})^{t-s-1}\right\rVert\left\lVert\bm{\Delta}_{s}\right\rVert\left\lVert\bm{\delta}_{s}\right\rVert\leq\eta\sum_{s=0}^{t-1}(1-\eta\lambda_{\min})^{t-s-1}\left\lVert\bm{\Delta}_{s}\right\rVert\left\lVert\bm{\delta}_{s}\right\rVert$ (34)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\rho\phi\cdot\eta\sum_{s=0}^{t-1}(1-\eta\lambda_{\min})^{t-s-1}\cdot\frac{3}{2}p_{s}\leq\frac{3}{2}\eta\rho\phi\cdot\sum_{s=0}^{t-1}(1-\eta\lambda_{\min})^{t-s-1}\cdot(1-\eta\lambda_{\min})^{s}p_{0}$
$=\frac{3}{2}\eta t\rho\phi\cdot p_{t-1}\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{1}{2}p_{t}.$
Here in $(a)$, we use $\left\lVert\bm{\Delta}_{t}\right\rVert\leq\rho\phi$ and the assumption $\left\lVert\bm{\delta}_{s}\right\rVert\leq p_{s}+q_{s}\leq\frac{3}{2}p_{s}$ for all $0\leq s\leq t-1$. In $(b)$, we use the fact that $t\leq\frac{2}{\eta\sqrt{\rho\epsilon}}\log\left(\frac{1}{\omega}\right)$ and $\phi=\frac{\sqrt{\epsilon}}{6\sqrt{\rho}}\log^{-1}\left(\frac{1}{\omega}\right)\leq\frac{1}{4}$, which implies
@@
Therefore, within $T=\log\left(\frac{4\phi}{\omega}\right)/\log\left(1+\eta\sqrt{\rho\epsilon}\right)$ iterations, we must have $\left\lVert\bm{\delta}_{T}\right\rVert\geq 2\phi$. Note that $\phi\leq\frac{1}{4}$ and $\log(1+\eta\sqrt{\rho\epsilon})\geq\frac{\eta\sqrt{\rho\epsilon}}{1+\eta\sqrt{\rho\epsilon}}\geq\frac{1}{2}\eta\sqrt{\rho\epsilon}$. Hence, we have $T\leq\frac{2}{\eta\sqrt{\rho\epsilon}}\log\left(\frac{1}{\omega}\right)$. This implies
@@
$\mathcal{X}_{\mathrm{stuck}}=\left\{\mathbf{x}_{0}\in\mathbb{B}_{\mathbf{z}}(\gamma):f(\mathbf{x}_{T})-f(\mathbf{z})\geq-F\text{ where }\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\nabla f(\mathbf{x}_{t}),\forall 0\leq t\leq T-1\right\}.$ (41)
@@
Using Lemma 4, we show that for any two GD sequences with initial points $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}\in\mathbb{B}_{\mathbf{z}}(\gamma)$ satisfying $\mathbf{x}_{0}-\mathbf{x}_{0}^{\prime}=\omega\mathbf{v}$, at least one of them lies outside the stuck region. We have
@@
$\mathbb{P}\left(\mathbf{x}_{0}\in\mathcal{X}_{\mathrm{stuck}}\right)=\frac{\operatorname{vol}\left(\mathcal{X}_{\text{stuck}}\right)}{\operatorname{vol}\left(\mathbb{B}^{d}(\gamma)\right)}\leq\frac{\omega\cdot\operatorname{vol}\left(\mathbb{B}^{d-1}(\gamma)\right)}{\operatorname{vol}\left(\mathbb{B}^{d}(\gamma)\right)}=\frac{\omega\Gamma(d/2+1)}{\sqrt{\pi}\gamma\Gamma(d/2+1/2)}$ $\leq\frac{\omega}{\gamma}\sqrt{\frac{d+1}{2\pi}}\leq\frac{\chi}{N_{\mathrm{perturb}}}.$ (44)
Here, $\operatorname{vol}(\mathcal{X})$ denotes the volume of the set $\mathcal{X}$. For the $d$-dimensional unit ball, we have $\operatorname{vol}(\mathbb{B}^{d}(1))=\frac{\pi^{d/2}}{\Gamma\left(\frac{d}{2}+1\right)}$, where $\Gamma(\cdot)$ is the Gamma function defined as $\Gamma\left(\frac{d}{2}+1\right)=\sqrt{\pi}\left(d-\frac{1}{2}\right)\cdot\left(d-\frac{3}{2}\right)\cdot\ldots\cdot\frac{1}{2}$. In the second inequality, we use $\frac{\Gamma(x+1)}{\Gamma(x+1/2)}<\sqrt{x+\frac{1}{2}}$ for all $x\geq 0$. ∎
@@ 4 Establishing Implicit Regularization of IPGD
Defining $\mathbf{u}^{\star}=\operatorname*{arg\,min}_{\mathbf{u}\in\mathcal{M}}\left\lVert\mathbf{u}-(\mathbf{x}+\mathbf{y})\right\rVert$ and $\mathbf{v}^{\star}=\operatorname*{arg\,min}_{\mathbf{v}\in\mathcal{M}}\left\lVert\mathbf{v}-\mathbf{x}\right\rVert$, we have:
@@
$\left\lVert\left(\mathbf{I}-\eta\nabla^{2}f\big(\bm{\mathbf{x}}^{\sharp}_{t}\big)\right)\mathbf{x}_{t}^{\perp}\right\rVert^{2}$ $=\left\lVert\sum_{i=1}^{d}(1-\eta\lambda_{i})\left\langle\mathbf{v}_{i},\bm{\mathbf{x}}^{\perp}_{t}\right\rangle\mathbf{v}_{i}\right\rVert^{2}=\sum_{i=1}^{d}(1-\eta\lambda_{i})^{2}\left\langle\mathbf{v}_{i},\bm{\mathbf{x}}^{\perp}_{t}\right\rangle^{2}$ (51)
@@
$\left\lVert\left(\mathbf{I}-\eta\nabla^{2}f\big(\bm{\mathbf{x}}^{\sharp}_{t}\big)\right)\mathbf{x}_{t}^{\perp}\right\rVert^{2}\leq\sum_{i:\lambda_{i}>0}\left(1-\eta\lambda_{i}\right)\left\langle\mathbf{v}_{i},\bm{\mathbf{x}}^{\perp}_{t}\right\rangle^{2}+\sum_{i:\lambda_{i}<0}\left(1-3\eta\lambda_{i}\right)\left\langle\mathbf{v}_{i},\bm{\mathbf{x}}^{\perp}_{t}\right\rangle^{2}$ (53)
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert^{2}+\eta\left(-\left\langle\bm{\mathbf{x}}^{\perp}_{t},\nabla^{2}_{+}f\big(\bm{\mathbf{x}}^{\sharp}_{t}\big)\bm{\mathbf{x}}^{\perp}_{t}\right\rangle-3\left\langle\bm{\mathbf{x}}^{\perp}_{t},\nabla^{2}_{-}f\big(\bm{\mathbf{x}}^{\sharp}_{t}\big)\bm{\mathbf{x}}^{\perp}_{t}\right\rangle\right)$
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert^{2}+\eta\left(r_{-}(\mathbf{x}_{t})+3r_{+}(\mathbf{x}_{t})\right)\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert^{2}\stackrel{{\scriptstyle(c)}}{{=}}(1+\eta r(\mathbf{x}_{t}))\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert^{2}.$
@@
$\geq\left(1+\frac{\eta}{2}R(\tau)+\frac{\eta\rho}{2}b_{t_{N_{\mathrm{perturb}}}-1}\right)b_{t_{N_{\mathrm{perturb}}}-1}+\gamma=b_{t_{N_{\mathrm{perturb}}}}+\gamma.$
@@
$\geq\left(1+\frac{\eta}{2}R(\tau)+\frac{\eta\rho}{2}\left(b_{t_{N_{\mathrm{perturb}}}}+\gamma\right)\right)\left(b_{t_{N_{\mathrm{perturb}}}}+\gamma\right)\geq b_{t_{N_{\mathrm{perturb}}}+1}.$
@@
Note that $a_{t+1}=a_{0}\prod_{s=0}^{t}\left(1+\eta R(\tau)+\eta\rho a_{s}\right)$. Hence, it suffices to prove $a_{t}\leq\frac{1}{\eta\rho T}$ for all $0\leq t\leq T-1$. We prove this by induction on $t$. It is easy to verify the base case $t=0$. Now, assume that $a_{s}\leq\frac{1}{\eta\rho T}$ for all $0\leq s\leq t$. Then, we have
[…]
@@
First, according to Lemmas 8 and 9, we have that for all $0\leq t\leq T$,
@@
where the final inequality uses $\left\lVert\mathbf{x}_{0}^{\perp}\right\rVert\leq\frac{\epsilon}{10L}$, $N_{\mathrm{perturb}}=\frac{C\sqrt{\rho}\Delta_{f}\left(\log^{3}\left(\frac{1}{\gamma}\right)+\log^{3}\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)}{\epsilon^{3/2}}$, $\gamma=O\left(\frac{\epsilon^{5/2}}{\sqrt{\rho}L\Delta_{f}}\log^{-3}\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)$, and $R(\tau)=O(\frac{1}{\eta T})$. Given $\left\lVert\mathbf{x}_{t}^{\perp}\right\rVert\leq\frac{\epsilon}{L}$ for all $0\leq t\leq T$, the gradient and Hessian of the objective function remain Lipschitz continuous along the entire trajectory $\{\mathbf{x}_{t}\}_{t=0}^{T}$. Thus, we can apply Theorem 2, which guarantees that, with probability at least $1-\chi$, IPGD outputs a point $\mathbf{x}_{T}$ that is an $\epsilon$-SOSP after $T=O\left(\frac{\Delta_{f}}{\eta\epsilon^{2}}\left(\log^{4}\left(\frac{1}{\gamma}\right)+\log^{4}\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)\right)$ iterations. Combining the above two results ensures that $\mathbf{x}_{T}$ is an $\epsilon$-$\mathcal{M}$-SOSP. ∎
@@ 5 Establishing Improved Convergence of IPGD with Additional Structures
$\left\lVert\mathbf{x}_{T}^{\perp}\right\rVert\leq\frac{\epsilon}{L},\quad\mathbf{x}_{T}\text{ is an }\bar{\epsilon}\text{-$\mathcal{M}$-SOSP}.$ (63)
@@
To control $(\textup{I})$, we write
$(\textup{I})^{2}$ $=\min_{\mathbf{x}^{\star}\in\mathcal{X}_{\mathcal{M}}}\left\lVert\mathbf{x}_{t}^{\sharp}-\eta\nabla f\big(\mathbf{x}_{t}^{\sharp}\big)-\mathbf{x}^{\star}\right\rVert^{2}\leq\left\lVert\mathbf{x}_{t}^{\sharp}-\eta\nabla f\big(\mathbf{x}_{t}^{\sharp}\big)-\mathcal{P}_{\mathcal{X}_{\mathcal{M}}}\big(\mathbf{x}_{t}^{\sharp}\big)\right\rVert^{2}$ (69)
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left(1-\frac{\eta\alpha}{2}\right)\left\lVert\mathbf{x}_{t}^{\sharp}-\mathcal{P}_{\mathcal{X}_{\mathcal{M}}}\big(\mathbf{x}_{t}^{\sharp}\big)\right\rVert^{2}+\left(\eta^{2}-\frac{\eta}{2\beta}\right)\left\lVert\nabla f\big(\mathbf{x}_{t}^{\sharp}\big)\right\rVert^{2}\stackrel{{\scriptstyle(b)}}{{\leq}}\left(1-\frac{\eta\alpha}{2}\right)\left\lVert\mathbf{x}_{t}^{\sharp}-\mathcal{P}_{\mathcal{X}_{\mathcal{M}}}\big(\mathbf{x}_{t}^{\sharp}\big)\right\rVert^{2},$
@@
In the last inequality, we use the triangle inequality and the basic inequality $\sqrt{1-x}\leq 1-\frac{1}{2}x$ for all $x\leq 1$. Next, we control $(\textup{II})$ as follows
$(\textup{II})$ $\leq\left\lVert\mathbf{x}_{t}^{\perp}\right\rVert+\eta\left\lVert\nabla f\big(\mathbf{x}_{t}^{\sharp}\big)-\nabla f(\mathbf{x}_{t})\right\rVert\leq\left(1+\eta L\right)\left\lVert\mathbf{x}_{t}^{\perp}\right\rVert.$ (71)
@@
Lemma 11 (Adapted from Lemma 47 in [ma2024convergence]).
@@
$\leq\left(1-\frac{\eta\alpha}{4}\right)\zeta+4\max_{0\leq t\leq T}\left\lVert\mathbf{x}_{t}^{\perp}\right\rVert\stackrel{{\scriptstyle(a)}}{{\leq}}\left(1-\frac{\eta\alpha}{4}\right)\zeta+O\left(\frac{\eta\alpha}{L}\epsilon\right)\stackrel{{\scriptstyle(b)}}{{\leq}}\zeta.$
In $(a)$, we use Equation 74. Step $(b)$ follows from the assumption that $\zeta\geq\frac{\epsilon}{2L}$. This completes the inductive step. Given that we established $\mathbf{x}^{\sharp}_{t}\in\mathcal{N}_{\mathcal{X}_{\mathcal{M}}}(\zeta),0\leq t\leq T$, we can invoke Lemma 10 for any $0\leq t\leq T-1$, which yields
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left(1-\frac{\eta\alpha}{4}\right)^{T}\cdot 2\zeta+\frac{8+4\eta L}{\eta\alpha}\max_{0\leq t\leq T}\left\{\left\lVert\mathbf{x}_{t}^{\perp}\right\rVert\right\}\stackrel{{\scriptstyle(b)}}{{\leq}}\left(1-\frac{\eta\alpha}{4}\right)^{T}\cdot 2\zeta+\frac{\epsilon}{2L}.$
@@
The proof follows by combining Proposition 2 and Proposition 3. ∎